# Kondo Screening and Indirect Magnetic Exchange through a Conventional Superconductor Studied by the Density-Matrix Renormalization Group

### v1

@@ I Introduction
Unconventional properties of physical systems are often a consequence of competing mechanisms at comparable energy scales. The Doniach diagram Doniach (1977) represents a prime example. Here, the individual Kondo screening Kondo (1964); Hewson (1993) of magnetic moments locally exchange-coupled to a Fermi sea competes with their mutual nonlocal magnetic interaction, which is mediated by the same Fermi sea. This Ruderman-Kittel-Kasuya-Yosida (RKKY) indirect exchange Ruderman and Kittel (1954); Kasuya (1956); Yosida (1957) $J_{\rm RKKY}\propto J^{2}$ is accessible by perturbation theory in the strength of the local exchange coupling $J$, while the Kondo effect is intrinsically non-perturbative with the Kondo temperature $T_{\rm K}\propto e^{-1/J}$ as the characteristic energy scale. The resulting Doniach diagram is a classical topic Jayaprakash et al. (1981); Jones and Varma (1987); Jones et al. (1988) of many-body theory and material science that has stimulated a continuous discussion and methodical advancements over several decades Fye et al. (1987); Sakai et al. (1990, 1993); Vojta et al. (2002); Zaránd et al. (2006); Jayatilaka et al. (2011); Titvinidze et al. (2012); Mitchell et al. (2012); Schwabe et al. (2013); Allerdt et al. (2015); Yevtushenko and Yudson (2018); Allerdt and Feiguin (2019); Zheng et al. (2020), see, e.g., Refs. Eickhoff et al. (2018); Eickhoff and Anders (2020) for an overview.
@@
The Kondo-vs.-RKKY competition is expected to be fundamentally different, if the electronic structure of the host material is gapped. For a single magnetic impurity in a finite system, i.e., for a “Kondo box” Thimm et al. (1999), and when the finite-size gap $\delta$ becomes comparable to the bulk Kondo temperature, the logarithmic Kondo correlations are truncated by the system size. Hence, for $\delta\gtrsim T^{\rm(bulk)}_{\rm K}$, the impurity spin cannot be Kondo screened because there are no low-energy excitations with energies below $T^{\rm(bulk)}_{\rm K}$ available. The physics resulting from the competition between $T_{\rm K}$ and $\delta$ has been studied extensively Schlottmann (2001, 2002); Simon and Affleck (2002); Cornaglia and Balseiro (2002); Franzese, G. et al. (2003); Rossi and Morr (2006); Hand et al. (2006); Bedrich et al. (2010).
@@
For several magnetic impurities Luo et al. (2005), i.e., for a finite “Kondo-vs.-RKKY box”, the truncated Kondo effect competes with an RKKY-like nonlocal exchange with an exponential distance dependence, reminiscent of the Bloembergen-Rowland-type Bloembergen and Rowland (1955). Particularly interesting is the “on resonance” case, where a singly occupied conduction-electron eigenstate right at the Fermi edge gives rise to a two-fold degenerate many-electron ground state and is thus available for the screening. It has been shown that this results in an unconventional reentrant competition at weak $J$ and small $\delta$ Schwabe et al. (2012). Later, this work has been generalized to an arbitrary number of magnetic moments in a metallic nanostructure Schwabe et al. (2015) and to the strong-$J$ limit Schwabe et al. (2014), where the low-energy physics is captured by generalized central-spin models. Ultrafast dynamical manipulations of the interplay between Kondo and RKKY exchange have been studied for small nanoring geometries Ydman et al. (2018).
@@
Several quantum spins coupled to a conventional s-wave superconductor represents another highly interesting class of systems with a competition between Kondo screening and indirect RKKY exchange. The competition is controlled by a gapped electronic structure and turns into a competition of a truncated “Kondo” effect (with exponentially decaying Kondo correlations) with a short-range exponentially suppressed indirect “RKKY” magnetic exchange. Within standard BCS theory, the gap size $2\Delta$ is related to the pairing strength $\Delta$ and is thus easily controlled via its temperature dependence. Opposed to Kondo-box systems, however, the ground state is a total-spin singlet at $J=0$, and thus there is no “on resonance” case. The ground-state properties are rather dominated by level crossings of subgap bound states Balatsky et al. (2006).
@@
Currently, there is a strong interest in chains of magnetic adatoms on superconductors Schneider et al. (2020); Kamlapure et al. (2021); von Oppen and Franke (2021); Friedrich et al. (2021); Deb et al. (2021); Mier et al. (2021); Liebhaber et al. (2022); Majek et al. (2022); Steiner et al. (2022); Žonda et al. (2023). The induced Yu-Shiba-Rusinov (YSR) bound states Yu (1965); Shiba (1968); Rusinov (1969) hybridize to form a subgap energy band that can host Majorana zero modes. As a platform for Majorana zero modes Nadj-Perge et al. (2013); Braunecker and Simon (2013); Pientka et al. (2013); Klinovaja et al. (2013) and topological superconductivity Pawlak et al. (2019); Flensberg et al. (2021) such systems have attracted much attention recently. The simplest system in this context consists of just two magnetic adatoms in proximity to a conventional superconductor Ruby et al. (2018); Choi et al. (2018); Kezilebieke et al. (2018); Beck et al. (2021). The importance of treating magnetic impurities or adatoms as quantum spins for the construction of the superconducting extension of Doniach’s phase diagram has been emphasized recently Schmid et al. (2022).
@@
In this paper, we study the case of two quantum spins with $s=\frac{1}{2}$ with antiferromagnetic exchange coupling $J$ to a conventional (BCS) superconductor with gap $2\Delta$. The two-impurity Kondo model poses a highly nontrivial many-body problem, which somewhat simplifies in case of a superconducting host, because the Kondo scale is cut by $\Delta$ and since the gap leads to exponentially decaying correlations. In the weak-$J$ regime the problem thus becomes accessible to standard perturbation theory. Still, an exact treatment requires state-of-the-art numerical methods, such as the numerical renormalization group (NRG) Wilson (1975); Krishna-murthy et al. (1980); Jones and Varma (1987); Bulla et al. (2008). The generic zero-temperature phase diagram has been studied using NRG Žitko et al. (2010, 2011); Yao et al. (2014) and has served as an important orientation for subsequent investigations Ruby et al. (2018); Friedrich et al. (2021); Kamlapure et al. (2021); von Oppen and Franke (2021); Majek et al. (2022); Steiner et al. (2022); Žonda et al. (2023). In particular, in a certain parameter regime, this exhibits a total-spin-$\frac{1}{2}$ “molecular doublet” phase, which is delocalized between the magnetic impurities, and other phases with dominating Kondo or RKKY correlations. However, within NRG, for technical reasons and related to the assumption of a flat density of states (see Ref. Yao et al. (2014), supplemental material), the inter-impurity indirect magnetic (RKKY) exchange $J_{\rm RKKY}$ is treated as an independent parameter.
@@
Here, we employ the density-matrix renormalization group (DMRG) White (1992); Schollwöck (2011) to study the ground-state phase diagram. The inter-impurity coupling is treated as an emergent interaction that depends on the details of the model and, in particular, on the distance $d$ between the impurity spins. On the other hand, the DMRG phase diagram is less generic and will depend on the details of the specific system geometry. The natural choice in the DMRG context is that of a one-dimensional chain geometry, and this is adopted here. Although this implies that the NRG and the DMRG results cannot be quantitatively mapped onto each other, it will be important to check, whether or not the general phase diagram topologies are the same. Due to the expected exponential decay of RKKY correlations, we focus on geometries with spin impurities locally exchange coupled to nearest and to next-nearest neighbor sites of the chain.
@@ II Model and symmetries
The Hamiltonian is invariant under SU(2) spin rotations and thus conserves the total spin ${\boldsymbol{S}}\equiv\sum_{i}{\boldsymbol{s}}_{i}+\sum_{m}{\boldsymbol{S}}_{m}$. At $\Delta=0$ and half-filling, there is an additional SU(2) symmetry in the charge sector (“hidden” SUc(2) charge symmetry Jones and Varma (1987)) generated by the components of the total pseudo-spin ${\boldsymbol{\eta}}\equiv\sum_{i}{\boldsymbol{\eta}}_{i}$, i.e., $[{\boldsymbol{\eta}},H]=0$. The $z$ component of the local pseudo-spin at site $i$ is defined as $\eta_{iz}=(c_{i\uparrow}^{\dagger}c_{i\uparrow}+c_{i\downarrow}^{\dagger}c_{i\downarrow}-1)/2$, and the $x$ and $y$ components, $\eta_{ix}=\frac{1}{2}(\eta_{i+}+\eta_{i-})$ and $\eta_{iy}=\frac{1}{2i}(\eta_{i+}-\eta_{i-})$, are given in terms of $\eta_{i+}=\epsilon_{i}c_{i\uparrow}^{\dagger}c_{i\downarrow}^{\dagger}$ and $\eta_{i-}=\epsilon_{i}c_{i\downarrow}c_{i\uparrow}$. Here, $\epsilon_{i}=(-1)^{i}$ is the usual sign factor for a bipartite lattice. The components of ${\boldsymbol{\eta}}_{i}$ span the charge $\mathfrak{su}(2)$ algebra with $[\eta_{i\alpha},\eta_{i\beta}]=\sum_{\gamma=x,y,z}\epsilon_{\alpha\beta\gamma}\eta_{i\gamma}$. We have $[s_{i\alpha},\eta_{i^{\prime}\alpha^{\prime}}]=0$ for arbitrary $\alpha,\alpha^{\prime}=x,y,z$.
@@
is proportional to the $x$ component of the total staggered pseudo-spin, the charge (pseudo-spin) SU(2) symmetry is broken. This implies that $\eta_{z}=(N-L)/2$ and hence $N$ are no longer conserved. Only the pseudo-spin $U(1)$ symmetry around the $x$ axis and the conservation of $\eta_{x}$ (“pseudo-charge”) with (half-)integer eigenvalues $q$ remains Yoshioka and Ohashi (2000). Furthermore, for $\Delta>0$ we still have conservation of the fermion parity $\Pi\equiv(-1)^{N}$. Spatial parity, unitarily represented by $P=\sum_{i\sigma}c_{i\sigma}^{\dagger}c_{L+1-i,\sigma}$ is conserved. For odd $L$ (only), $P$ additionally commutes with $\eta_{x}$.
@@ III DMRG computations
To compute ground-state expectation values and correlations, we use a variational approach to optimize matrix-product states. Our algorithm is based on the ideas of DMRG White (1992); Schollwöck (2011). It relies on the finding that the ground state of gapped one-dimensional Hamiltonians with short-range hopping and interactions obeys an entropic area law, i.e., the entanglement entropy does not scale with the system size. Matrix-product states provide a well-suited platform for the handling of such low-entanglement quantum states. Furthermore, our code fully exploits the non-abelian spin-SU(2) symmetry by using symmetry-reduced basis states McCulloch and Gulásci (2002); Weichselbaum (2012), and by using a lossless compression of matrix-product operators Hubig et al. (2017). The bond dimension of symmetry-reduced state matrices must take values of the order of $10^{3}$ to provide a numerically precise approximation of the quantum states under study. With increasing $\Delta$, the required bond dimension becomes smaller, because the ground-state entanglement decreases when the correlations become predominantly local. The convergence of the algorithm is controlled by the measure $\Delta E^{2}/L=(\langle H^{2}\rangle-\langle H\rangle^{2})/L$ and is additionally monitored by the change of the eigenvalues of the state matrices during the local optimization. The implementation of the spin-SU(2) symmetry allows us to target spin quantum numbers without numerical interference of energetically close states from other subspaces. This is important in order to distinguish between the ground-state energies for different spin quantum numbers with a high degree of accuracy.
@@ IV.1 Phase diagram for distance $d=1$
Figure 1: Impurity spin correlation $\left\langle\boldsymbol{S}_{1}\boldsymbol{S}_{2}\right\rangle$ (see color code) as obtained by DMRG for a system consisting of two $s=\frac{1}{2}$ impurity spins ($M=2$), exchange coupled to the nearest-neighbor ($d=1$) central sites of a one-dimensional chain with $L=82$ sites as function of $J$ and $\Delta$. The nearest-neighbor hopping $t=1$ fixes the energy scale.
The competition between BCS-singlet formation, Kondo screening, and the emergent indirect RKKY exchange crucially depends on the distance $d$ between the impurity spins. For nearest neighbors, $d=1$, Fig. 1 shows the impurity-spin correlation $\langle{\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\rangle$ in the ground state of a long chain ($L=82$ sites). The impurity spins are coupled to sites in the center of the chain.
@@
However, the character of the singlet ground state is very different in the different limits. For $J\gg\Delta,t$ the ground state exhibits strong antiferromagnetic Kondo correlations $\langle{\boldsymbol{S}}_{m}{\boldsymbol{s}}_{i_{m}}\rangle\approx-\frac{3}{4}$ ($m=1,2$) at the impurity sites. For $J\to\infty$, these essentially decouple the left and the right remainder of chain, $i<i_{1}$ and $i>i_{2}$, which then form two individual conventional BCS singlets.
Our DMRG results are almost perfectly independent of the system size as long as $\Delta$ is larger than the finite-size gap $\delta\sim 1/L$. We found that this limits our studies of ground-state correlations to $\Delta\gtrsim 0.1$. Given this limitation, our results for $L=82$ are fully converged. This is also true for all DMRG results throughout the paper.

### v2

@@ I Introduction
Unconventional properties of physical systems are often a consequence of competing mechanisms at comparable energy scales. The Doniach diagram [1] represents a prime example. Here, the individual Kondo screening [2, 3] of magnetic moments locally exchange-coupled to a Fermi sea competes with their mutual nonlocal magnetic interaction, which is mediated by the same Fermi sea. This Ruderman-Kittel-Kasuya-Yosida (RKKY) indirect exchange [4, 5, 6] $J_{\rm RKKY}\propto J^{2}$ is accessible by perturbation theory in the strength of the local exchange coupling $J$, while the Kondo effect is intrinsically non-perturbative with the Kondo temperature $T_{\rm K}\propto e^{-1/J}$ as the characteristic energy scale. The resulting Doniach diagram is a classical topic [7, 8, 9] of many-body theory and material science that has stimulated a continuous discussion and methodical advancements over several decades [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], see, e.g., Refs. [23, 24] for an overview.
@@
The Kondo-vs.-RKKY competition is expected to be fundamentally different, if the electronic structure of the host material is gapped. For a single magnetic impurity in a finite system, i.e., for a “Kondo box” [25], and when the finite-size gap $\delta$ becomes comparable to the bulk Kondo temperature, the logarithmic Kondo correlations are truncated by the system size. Hence, for $\delta\gtrsim T^{\rm(bulk)}_{\rm K}$, the impurity spin cannot be Kondo screened because there are no low-energy excitations with energies below $T^{\rm(bulk)}_{\rm K}$ available. The physics resulting from the competition between $T_{\rm K}$ and $\delta$ has been studied extensively [26, 27, 28, 29, 30, 31, 32, 33].
@@
For several magnetic impurities [34], i.e., for a finite “Kondo-vs.-RKKY box”, the truncated Kondo effect competes with an RKKY-like nonlocal exchange with an exponential distance dependence at large distances, reminiscent of the Bloembergen-Rowland-type [35]. Particularly interesting is the “on resonance” case, where a singly occupied conduction-electron eigenstate right at the Fermi edge gives rise to a two-fold degenerate many-electron ground state and is thus available for the screening. It has been shown that this results in an unconventional reentrant competition at weak $J$ and small $\delta$ [36]. Later, this work has been generalized to an arbitrary number of magnetic moments in a metallic nanostructure [37] and to the strong-$J$ limit [38], where the low-energy physics is captured by generalized central-spin models. Ultrafast dynamical manipulations of the interplay between Kondo and RKKY exchange have been studied for small nanoring geometries [39].
@@
Several quantum spins coupled to a conventional s-wave superconductor represents another highly interesting class of systems with a competition between Kondo screening and indirect RKKY exchange. The competition is controlled by a gapped electronic structure and turns into a competition of a truncated “Kondo” effect (with exponentially decaying Kondo correlations) with an attenuated indirect “RKKY” magnetic exchange, which is exponentially suppressed at large distances. Within standard BCS theory, the gap size $2\Delta$ is related to the pairing strength $\Delta$ and is thus easily controlled via its temperature dependence. Opposed to Kondo-box systems, however, the ground state is a total-spin singlet at $J=0$, and thus there is no “on resonance” case. The ground-state properties are rather dominated by level crossings of subgap bound states [40].
@@
Currently, there is a strong interest in chains of magnetic adatoms on superconductors [41, 42, 43, 44, 45, 46, 47, 48, 49, 50]. The induced Yu-Shiba-Rusinov (YSR) bound states [51, 52, 53] hybridize to form a subgap energy band that can host Majorana zero modes. As a platform for Majorana zero modes [54, 55, 56, 57] and topological superconductivity [58, 59] such systems have attracted much attention recently. The simplest system in this context consists of just two magnetic adatoms in proximity to a conventional superconductor [60, 61, 62, 63]. The importance of treating magnetic impurities or adatoms as quantum spins for the construction of the superconducting extension of Doniach’s phase diagram has been emphasized recently [64].
@@
In this paper, we study the case of two quantum spins with $s=\frac{1}{2}$ with antiferromagnetic exchange coupling $J$ to a conventional (BCS) superconductor with gap $2\Delta$. The two-impurity Kondo model poses a highly nontrivial many-body problem, which somewhat simplifies in case of a superconducting host, because the Kondo scale is cut by $\Delta$ and since the gap leads to exponentially decaying correlations. In the weak-$J$ regime the problem thus becomes accessible to standard perturbation theory. Still, an exact treatment requires state-of-the-art numerical methods, such as the numerical renormalization group (NRG) [65, 66, 8, 67]. The generic zero-temperature phase diagram has been studied using NRG [68, 69, 70] and has served as an important orientation for subsequent investigations [60, 44, 42, 43, 48, 49, 50]. In particular, in a certain parameter regime, this exhibits a total-spin-$\frac{1}{2}$ “molecular doublet” phase, which is delocalized between the magnetic impurities, and other phases with dominating Kondo or RKKY correlations. However, within NRG, for technical reasons and related to the assumption of a flat density of states (see Ref. [70], supplemental material), the inter-impurity indirect magnetic (RKKY) exchange $J_{\rm RKKY}$ is treated as an independent parameter.
@@
Here, we employ the density-matrix renormalization group (DMRG) [71, 72] to study the ground-state phase diagram. The inter-impurity coupling is treated as an emergent interaction that depends on the details of the model and, in particular, on the distance $d$ between the impurity spins. On the other hand, the DMRG phase diagram is less generic and will depend on the details of the specific system geometry. The natural choice in the DMRG context is that of a one-dimensional chain geometry, and this is adopted here. Although this implies that the NRG and the DMRG results cannot be quantitatively mapped onto each other, it will be important to check, whether or not the general phase diagram topologies are the same. Due to the expected exponential decay of RKKY correlations at large distances, we focus on geometries with spin impurities locally exchange coupled to nearest and to next-nearest neighbor sites of the chain.
@@ II Model and symmetries
The Hamiltonian is invariant under SU(2) spin rotations and thus conserves the total spin ${\boldsymbol{S}}\equiv\sum_{i}{\boldsymbol{s}}_{i}+\sum_{m}{\boldsymbol{S}}_{m}$. At $\Delta=0$ and half-filling, there is an additional SU(2) symmetry in the charge sector (“hidden” SUc(2) charge symmetry [8]) generated by the components of the total pseudo-spin ${\boldsymbol{\eta}}\equiv\sum_{i}{\boldsymbol{\eta}}_{i}$, i.e., $[{\boldsymbol{\eta}},H]=0$. The $z$ component of the local pseudo-spin at site $i$ is defined as $\eta_{iz}=(c_{i\uparrow}^{\dagger}c_{i\uparrow}+c_{i\downarrow}^{\dagger}c_{i\downarrow}-1)/2$, and the $x$ and $y$ components, $\eta_{ix}=\frac{1}{2}(\eta_{i+}+\eta_{i-})$ and $\eta_{iy}=\frac{1}{2i}(\eta_{i+}-\eta_{i-})$, are given in terms of $\eta_{i+}=\epsilon_{i}c_{i\uparrow}^{\dagger}c_{i\downarrow}^{\dagger}$ and $\eta_{i-}=\epsilon_{i}c_{i\downarrow}c_{i\uparrow}$. Here, $\epsilon_{i}=(-1)^{i}$ is the usual sign factor for a bipartite lattice. The components of ${\boldsymbol{\eta}}_{i}$ span the charge $\mathfrak{su}(2)$ algebra with $[\eta_{i\alpha},\eta_{i\beta}]=\sum_{\gamma=x,y,z}\epsilon_{\alpha\beta\gamma}\eta_{i\gamma}$. We have $[s_{i\alpha},\eta_{i^{\prime}\alpha^{\prime}}]=0$ for arbitrary $\alpha,\alpha^{\prime}=x,y,z$.
@@
is proportional to the $x$ component of the total staggered pseudo-spin, the charge (pseudo-spin) SU(2) symmetry is broken. This implies that $\eta_{z}=(N-L)/2$ and hence $N$ are no longer conserved. Only the pseudo-spin $U(1)$ symmetry around the $x$ axis and the conservation of $\eta_{x}$ (“pseudo-charge”) with (half-)integer eigenvalues $q$ remains [73]. Furthermore, for $\Delta>0$ we still have conservation of the fermion parity $\Pi\equiv(-1)^{N}$. Spatial parity, unitarily represented by $P=\sum_{i\sigma}c_{i\sigma}^{\dagger}c_{L+1-i,\sigma}$ is conserved. For odd $L$ (only), $P$ additionally commutes with $\eta_{x}$.
@@ III DMRG computations
To compute ground-state expectation values and correlations, we use a variational approach to optimize matrix-product states. Our algorithm is based on the ideas of DMRG [71, 72]. It relies on the finding that the ground state of gapped one-dimensional Hamiltonians with short-range hopping and interactions obeys an entropic area law, i.e., the entanglement entropy does not scale with the system size. Matrix-product states provide a well-suited platform for the handling of such low-entanglement quantum states. Furthermore, our code fully exploits the non-abelian spin-SU(2) symmetry by using symmetry-reduced basis states [74, 75], and by using a lossless compression of matrix-product operators [76]. The bond dimension of symmetry-reduced state matrices must take values of the order of $10^{3}$ to provide a numerically precise approximation of the quantum states under study. With increasing $\Delta$, the required bond dimension becomes smaller, because the ground-state entanglement decreases when the correlations become predominantly local. The convergence of the algorithm is controlled by the measure $\Delta E^{2}/L=(\langle H^{2}\rangle-\langle H\rangle^{2})/L$ and is additionally monitored by the change of the eigenvalues of the state matrices during the local optimization. For all results presented, we have achieved a total-energy variance per site smaller than $\Delta E^{2}/L=10^{-9}$. The implementation of the spin-SU(2) symmetry allows us to target spin quantum numbers without numerical interference of energetically close states from other subspaces. This is important in order to distinguish between the ground-state energies for different spin quantum numbers with a high degree of accuracy.
@@ IV.1 Phase diagram for distance $d=1$
Figure 1: Impurity spin correlation $\left\langle\boldsymbol{S}_{1}\boldsymbol{S}_{2}\right\rangle$ (see color code, upper plot) and Kondo correlation function $\langle{\boldsymbol{S}}_{m}{\boldsymbol{s}}_{m}\rangle$ (lower plot) as obtained by DMRG for a system consisting of two $s=\frac{1}{2}$ impurity spins ($M=2$), exchange coupled to the nearest-neighbor ($d=1$) central sites of a one-dimensional chain with $L=82$ sites as function of $J$ and $\Delta$. The nearest-neighbor hopping $t=1$ fixes the energy scale.
The competition between BCS-singlet formation, Kondo screening, and the emergent indirect RKKY exchange crucially depends on the distance $d$ between the impurity spins. For nearest neighbors, $d=1$, the upper plot in Fig. 1 shows the impurity-spin correlation $\langle{\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\rangle$ in the ground state of a long chain ($L=82$ sites). The impurity spins are coupled to sites in the center of the chain.
@@
However, the character of the singlet ground state is very different in the different limits. For $J\gg\Delta,t$ the ground state exhibits strong antiferromagnetic Kondo correlations $\langle{\boldsymbol{S}}_{m}{\boldsymbol{s}}_{i_{m}}\rangle\approx-\frac{3}{4}$ ($m=1,2$) at the impurity sites, see lower plot in Fig. 1. For $J\to\infty$, these essentially decouple the left and the right remainder of chain, $i<i_{1}$ and $i>i_{2}$, which then form two individual conventional BCS singlets. Generally, the local Kondo correlations are strongest, $\langle{\boldsymbol{S}}_{m}{\boldsymbol{s}}_{i_{m}}\rangle\to-\frac{3}{4}$, for strong $J$ and small $\Delta$. Contrary, for weak $J$ and large $\Delta$, we have $\langle{\boldsymbol{S}}_{m}{\boldsymbol{s}}_{i_{m}}\rangle\to 0$. This is just complementary to the impurity-spin correlation $\langle{\boldsymbol{S}}_{1}{\boldsymbol{S}}_{2}\rangle$. The transition between these limits is featureless and smooth.
Our DMRG results are almost perfectly independent of the system size as long as $\Delta$ is larger than the finite-size gap $\delta\sim 1/L$. We found that this limits our studies of ground-state correlations to $\Delta\gtrsim 0.1$. Given this limitation, our results for $L=82$ are fully converged. This is also true for all DMRG results throughout the paper. We also note that the results for the correlation functions and, at finite $\Delta$, the total spin $S$ are independent of whether $L$ is selected as even or odd (while keeping $\langle N\rangle=L$). The same holds for the case $d=2$ discussed below.